Please help me solve this algebra/geometry question I haven't been able to come up with anything and frankly I am completely confused as to what is going on in this problem.
In a parallelogram, a base and a corresponding height are in the ratio 5 : 2. The area is 250 cm2. Find the lengths of the base and the corresponding height. (Hint: Use 2x for the base and 5x for the height of the parallelogram.)

Answer by ewatrrr(24785)


Hi
Area(parallelogram) = b*h = 2x*5x = 250cm^2
10x^2 = 250 cm^2
x^2 = 25
x = 5cm (tossing out negative solution for a measure)
Base is 10cm and Height is 25cm
CHECKING our Answer***
10cm*25cm = 250 cm^2